OpenStudy (anonymous):

Write an equation for the line tangent to the circle x2 + y2 = 17, at the point (4, 1). Enter your answer in point-slope form: y - y1 = m(x - x1); example: y+3=-5(x-2)

OpenStudy (anonymous):

we know radius and tangent are at rt angles slope of radius passing through (0,0) and (4,1) =(1-0)/(4-0)=1/4 slope of tangent is -1/(1/4)=-4 hence eq of tangent passing through (4,1) (y-1)=-4(x-4)

OpenStudy (anonymous):

further 4x+1=17 on re arrnging the above eq.